# Faster-than-Nyquist Signaling in the Finite Time-Bandwidth Product Regime

### v3

@@ II-C $N$-parallel Gaussian channel formulation for FTN
Here, $\tilde{\mathbf{x}}\triangleq\mathbf{U}^{T}\mathbf{x}$ is the transformed input vector, which remains u.i.d. with unit variance due to $\text{cov}(\tilde{\mathbf{x}})=\text{cov}(\mathbf{x})$. [^4] The u.i.d. assumption, $\text{cov}(\tilde{\mathbf{x}})=\mathbf{I}$, means that the input must satisfy the average power constraint, $\mathbb{E}\{\lvert\tilde{x}_{n}\rvert^{2}\}=1$ for all $n$. In the MCCR analysis, sometimes it will be convenient to impose a stricter condition of constant input power for all realizations, i.e., $\lvert\tilde{x}_{n}\rvert^{2}=1$ for all $n$. The constant power implies that $\frac{1}{N}\|\mathbf{x}\|^{2}=1$ (i.e., equal power for individual codewords) due to $\|\mathbf{x}\|=\|\tilde{\mathbf{x}}\|$. We show in section III-C that the inputs $\{\tilde{x}_{n}\}$ under the constant power constraint can achieve the channel capacity as TBP tends to infinity (see also [scarlett2016dispersion] for asymptotic optimality of constant input power). The type of constraint used will be stated explicitly in the MCCR analysis. The vector $\tilde{\mathbf{z}}\triangleq\sqrt{\frac{N_{0}}{PT_{x}/N}}\Lambda^{-1}\mathbf{U}^{T}\mathbf{z}$ contains independent Gaussian noise samples, where the noise variance for the $n$-th subchannel is $\sigma_{n}^{2}=(\rho\frac{\Omega}{N}\lambda_{n})^{-1}$, and $\rho\triangleq\frac{P}{2WN_{0}}$ is the signal-to-noise ratio (SNR). The $n$-th subchannel therefore has the signal-to-noise ratio of $\text{SNR}_{n}=1/\sigma_{n}^{2}=\rho\frac{\Omega}{N}\lambda_{n}$. We note that FTN can increase the number of parallel channels by lowering $\tau$ in a fixed TBP.
@@
where the folded-spectrum is defined by $\hat{p}_{\text{folded}}(f)\triangleq\sum_{k=-\infty}^{\infty}\left|\hat{p}\left(f-\frac{k}{\tau T}\right)\right|^{2}$. Closed-form expressions for the folded-spectrum are available for many pulses of interest [kim2016properties]. While (12) yields an excellent approximation in many cases, its accuracy is not guaranteed when TBP is low, which is the focus of this work. For this reason, all results presented in this work are based on the direct numerical computation of the eigenvalues.
@@ III-A Capacity of FTN signaling
$\boxed{C_{\text{FTN}}=\frac{1}{2W}\int_{-\frac{1}{2\tau T}}^{\frac{1}{2\tau T}}log_{2}(1+\rho 2W\hat{p}_{\text{folded}}(f))df.}$ (13)
@@ V-A Signaling rate
One of the key benefits of FTN signaling is its ability to increase the number of parallel channels for a given TBP. To utilize the maximum number of signaling dimensions, the FTN time-acceleration factor $\tau$ can be set to yield $N^{*}$ channels as follows:
@@
We see from (20) that $\tau^{*}$ approaches $\tau_{0}$ in the asymptotic TBP regime (i.e., as $\Omega\to\infty$). In the finite TBP regime, however, $\eta$ is typically smaller than $c$ for practical ranges of $\epsilon_{W}$, which implies $\tau^{*}<\tau_{0}$ is required to utilize all $N^{*}$ dimensions. [^7] The minimum $c$ needed to meet a given OOB power constraint can be inferred from Fig. 4, which plots the OOB energy versus $c$ for various pulses. For instance, $c\geq 10.52$ for the RRC pulse with $\beta=0.5$ when $\epsilon_{W}=10^{-4}$. Fig. 5 plots the normalized SNRs of the parallel channels ($\sigma_{n}^{-2}/\rho$) for FTN systems with various $\tau$ values. As shown in solid lines, using $\tau_{0}$ activates only a partial number of channels, especially in low-TBP regimes, indicating that there is a benefit of selecting $\tau<\tau_{0}$ in the finite TBP regime. Setting $\tau=\tau^{*}$ yields $N^{*}$ channels as also indicated in the figure. The levels of these additional channels scale linearly with $\rho$ and thus the rate gains are higher in the high SNR. There is a diminishing return on rate gains for $\tau<\tau^{*}$, as the additional channels beyond $N^{*}$ possess diminishing SNRs and thus contribute little to the overall rate.
@@ V-B Time-limited pulse design
Numerous pulse designs have been proposed for FTN signaling. Jaffal and Alvarado [jaffal2022pulses] considered time-limited pulses based on a weighted combination of normalized and time-truncated PSWFs. The weights were optimized to minimize residual ISI power—specifically, ISI beyond the memory of the equalizer—while meeting an OOB constraint. Jaffal and Abou-Faycal [jaffal2019achievable] used a similar linear combination of PSWFs but with the objective of maximizing the capacity of a bandlimited channel. Anderson [anderson2017bandwidth, Ch. 7] explored several optimized pulses, including a modified principal PSWF, that simultaneously satisfy OOB and OOI requirements; however, these idealized pulses have infinite time and frequency support. Makarov et al. [makarov2020optimizing] proposed a time-limited pulse based on a Fourier-series (FS) expansion, optimized to minimize OOB power subject to a specific autocorrelation constraint. Despite these advanced designs, the time-truncated RRC pulse remains the common choice in the FTN literature due to its versatility and wide adoption in communication standards.
@@
We now compare the performance of the optimized pulse against other common pulses. Fig. 6 plots $R_{\text{NA}}$ for FTN systems using various pulses, alongside the PSWF benchmarks as detailed in Appendix E. The considered pulses include our optimized pulse from Table I (labeled ”modified Makarov”), time-truncated RRC pulses with various $\beta$, Anderson’s modified PSWF [anderson2017bandwidth], the principal PSWF pulse $\phi_{c,0}(t)$ (labeled ”PSWF0”), and Gaussian pulse. [^8] Anderson’s pulse is given by $p(t)=\alpha_{1}\psi_{c,0}(t)\text{ for }|t|\leq T/2$ and $p(t)=\alpha_{2}\psi_{c,0}(t)\text{ for }|t|>T/2$, where $\alpha_{1}$ and $\alpha_{2}$ are chosen so that OOB and OOI are both equal to $\epsilon$. See [anderson2017bandwidth] for details. Anderson’s pulse is further time-truncated to $|t|<T_{p}/2$ and energy normalized in our work. Gaussian (Gauss) is a time-truncated and energy normalized Gaussian function with the parameter $\sigma=Q^{-1}\left(\epsilon_{W}/2\right)/(2\sqrt{2}\pi W)$, chosen to make OOB equal to $\epsilon_{W}$. For each pulse, the minimum pulse-width (or the minimum $c=2WT_{p}$) needed to meet the OOB constraint $\epsilon_{W}=10^{-4}$ is obtained from a plot of OOB energy versus $c$ in Fig. 4. The Nyquist symbol times are given by $T=\frac{1+\beta}{2W}$ for RRC, $T=T_{p}$ for the principal PSWF and Gaussian. $T$ is numerically computed for Anderson so that the OOB constraint is met.
@@ V-C Coding design
The FTN receiver implements three-stage turbo equalization between the FTN channel equalizer, URC decoder, and RSC decoder that iteratively exchange the extrinsic information of the coded bits. Instead of the low-complexity frequency-domain FTN equalizer considered in [sugiura2015frequency], we use the optimal maximum a posteriori (MAP) equalizer with $L$-memory that is operating on the trellis description of the FTN channel model (9). Leveraging on the fact that the FTN channel model is an instance of the Ungerboeck observation model, the MAP equalizer is implemented using the BCJR algorithm with modified branch metrics (see eq. (11) in [rusek2015ungerboeck]) [^9] The computational complexity of the optimal MAP equalizer is $\mathcal{O}(M^{L}N)$ for $L$-memory equalizer with $M$ modulation order and $N$ symbols. The exponential complexity can be replaced by polynomial (or even linear) complexity using reduced-complexity FTN equalizers at some performance loss [sugiura2015frequency], [bedeer2017very]. Fortunately, $L$ generally need not be large, since RRC pulses with $\beta$ close to 1 that is shown to be near-optimal in low TBP has fast temporal roll off and generates short ISI even when $\tau\ll 1$. As TBP increases, larger $\beta$ is needed but at the same time $\tau$ need not be small (both $\tau^{*}$ and $\tau_{0}$ get closer to 1) thus ISI remains short. Channel shortening technique [fan2018mlse] can be used to further lower $L$ if needed. Numerical simulations confirmed that $L=3$ was sufficient to perform near MCCR in most cases of interest.. The MAP equalizer and the URC decoder form an inner turbo loop and exchange the extrinsic information about the URC encoded bits. The URC decoder generates the extrinsic information of both the URC encoded bits and the RSC encoded bits. After the inner loop iterates for $I_{\text{in}}$ times, the outer turbo loop between the URC decoder and the RSC decoder is run, in which the extrinsic information about the RSC encoded bits is exchanged. The inner loop then restarts with the updated extrinsic information from the RSC decoder becoming the a prior information about the URC encoded bits. The outer loop iterates for $I_{\text{out}}$, with which the total number of turbo iteration is $I_{\text{in}}I_{\text{out}}$.
@@ VI Conclusion
From a design perspective, we identified the optimal time-acceleration factor needed to maximize the available signaling dimensions and showed that with optimized pulse shaping, FTN’s performance can closely approach the theoretical benchmark set by PSWFs. Furthermore, we validated these theoretical gains with a practical three-stage turbo-equalization coding scheme that performs close to the MCCR bound, even for a TBP as low as $\Omega=132$.
Collectively, these results confirm that FTN effectively reduces the performance penalty from finite-length coding by creating more signaling dimensions within a constrained TBP. The fundamental benefits identified are expected to extend to more complex environments, such as fading, MIMO, and multi-user channels, positioning FTN as a valuable technique for future communication systems.
@@ Appendix A Proof of Theorem 2
$i_{M}(\tilde{x}_{n};\tilde{y}_{n})\triangleq\ln\left(\frac{p_{\tilde{y}|\tilde{x}}(\tilde{y}_{n}|\tilde{x}_{n})}{q_{\tilde{y}}(\tilde{y}_{n})}\right).$
@@
$i_{M}(\tilde{x}_{n};\tilde{y}_{n})=\ln\left(1+\frac{1}{\sigma_{n}^{2}}\right)-\frac{|\tilde{y}_{n}-(1+\sigma_{n}^{2})\tilde{x}_{n}|^{2}}{\sigma_{n}^{2}(1+\sigma_{n}^{2})}+|\tilde{x}_{n}|^{2}.$ (21)
@@ Appendix B Asymptotic expansion of the MC bound and proof of Proposition 1
$P_{\text{MD}}(\gamma)=P\left[\sum_{n=0}^{N-1}i_{M}(\tilde{x}_{n};\tilde{y}_{n})\leq N\gamma\middle|\text{H}_{1},\tilde{\mathbf{x}}\right$],
where $i_{M}(\tilde{x}_{n};\tilde{y}_{n})$ is the mismatched information density and given by (21) for the FTN channel. We will need its first three moments, which are given in the lemma below.
@@
(moments of mismatched information density): Let $|\tilde{x}_{n}|=1$. The mean and the variance of the mismatched information density $i_{M}(\tilde{x}_{n};\tilde{y}_{n})$ under $\tilde{y}_{n}\sim p_{\tilde{y}\mid\tilde{x}}=\mathcal{CN}(\tilde{x}_{n},\sigma_{n}^{2})$ and for a fixed $\tilde{x}_{n}$ are
$\begin{split}&c_{n}\triangleq\mathbb{E}_{\tilde{y}_{n}}[i_{M}(\tilde{x}_{n};\tilde{y}_{n})]=\ln\left(1+\tfrac{1}{\sigma_{n}^{2}}\right),\\ &v_{n}\triangleq\text{var}_{\tilde{y}_{n}}[i_{M}(\tilde{x}_{n};\tilde{y}_{n})]=1-\frac{1}{(1+\frac{1}{\sigma_{n}^{2}})^{2}}.\end{split}$
The third absolute central moment, $\theta_{n}\triangleq\mathbb{E}_{\tilde{y}_{n}}[|i_{M}(\tilde{x}_{n};\tilde{y}_{n})-c_{n}|^{3}]$, can be bounded as
@@
$(\mathbb{E}_{\tilde{y}_{n}}[|i_{M}(\tilde{x}_{n};\tilde{y}_{n})-c_{n}|^{2}])^{\frac{1}{2}}\leq\theta_{n}^{\frac{1}{3}}\leq(\mathbb{E}_{\tilde{y}_{n}}[|i_{M}(\tilde{x}_{n};\tilde{y}_{n})-c_{n}|^{4}])^{\frac{1}{4}}.$

### v4

@@ II-C $N$-parallel Gaussian channel formulation for FTN
Here, $\tilde{\mathbf{x}}\triangleq\mathbf{U}^{T}\mathbf{x}$ is the transformed input vector, which remains u.i.d. with unit variance due to $\text{cov}(\tilde{\mathbf{x}})=\text{cov}(\mathbf{x})$. [^4] The u.i.d. assumption, $\text{cov}(\tilde{\mathbf{x}})=\mathbf{I}$, means that the input must satisfy the average power constraint, $\mathbb{E}\{\lvert\tilde{x}_{n}\rvert^{2}\}=1$ for all $n$. In the MCCR analysis, sometimes it will be convenient to impose a stricter condition of constant input power for all realizations, i.e., $\lvert\tilde{x}_{n}\rvert^{2}=1$ for all $n$. The constant power implies that $\frac{1}{N}\|\mathbf{x}\|^{2}=1$ (i.e., equal power for individual codewords) due to $\|\mathbf{x}\|=\|\tilde{\mathbf{x}}\|$. We show in section III-C that the inputs $\{\tilde{x}_{n}\}$ under the constant power constraint can achieve the channel capacity as TBP tends to infinity (see also [scarlett2016dispersion] for asymptotic optimality of constant input power). The type of constraint used will be stated explicitly in the MCCR analysis. The vector $\tilde{\mathbf{z}}\triangleq\sqrt{\frac{N_{0}}{PT_{x}/N}}\mathbf{\Lambda}^{-1}\mathbf{U}^{T}\mathbf{z}$ contains independent Gaussian noise samples, where the noise variance for the $n$-th subchannel is $\sigma_{n}^{2}=(\rho\frac{\Omega}{N}\lambda_{n})^{-1}$, and $\rho\triangleq\frac{P}{2WN_{0}}$ is the signal-to-noise ratio (SNR). The $n$-th subchannel therefore has the signal-to-noise ratio of $\text{SNR}_{n}=1/\sigma_{n}^{2}=\rho\frac{\Omega}{N}\lambda_{n}$. We note that FTN can increase the number of parallel channels by lowering $\tau$ in a fixed TBP.
@@
where the folded-spectrum is defined by $\hat{p}_{\text{folded}}(f)\triangleq\sum_{k=-\infty}^{\infty}\left|\hat{p}\left(f-\frac{k}{\tau T}\right)\right|^{2}$. Closed-form expressions for the folded-spectrum are available for many pulses of interest [kim2016properties].
Figure 2: Normalized SNRs of the parallel channels ($\sigma_{n}^{-2}/\rho$) over normalized channel indices $n/\Omega$ for $n=0,1,\cdots,N-1$ for varying OOB levels and TBPs. RRC pulses with roll-off $\beta=0.5$, time-truncated so that OOB energy satisfies $\epsilon_{W}=10^{-2}$ or $10^{-4}$, and $\tau=1/(1+\beta)$ considered.
Fig. 2 plots the normalized SNRs of the $N$-parallel channels (i.e., $\sigma_{n}^{-2}/\rho=\tfrac{\Omega}{N}\lambda_{n}$) for varying OOB levels and TBPs. These are plotted over normalized channel indices $\frac{n}{\Omega}$ for $n=0,1,\cdots,N-1$. The eigenvalues $\{\lambda_{n}\}$ are computed numerically, while the folded-spectrum (12) is used for $\Omega\to\infty$. We observe that as TBP increases the normalized SNRs converge towards the asymptote (12), but the convergence is slow for more stringent OOB levels (e.g., $\epsilon_{W}=10^{-4}$). These imply that while (12) yields an excellent approximation in many cases, its accuracy is not guaranteed when TBP is low, which is the focus of this work. Deriving exact analytical eigenvalues of $H$ in the finite TBP regime is a challenging mathematical problem. For these reasons, all results presented in this work are based on the direct numerical computation of the eigenvalues.
@@ III-A Capacity of FTN signaling
$\boxed{C_{\text{FTN}}=\frac{1}{2W}\int_{-\frac{1}{2\tau T}}^{\frac{1}{2\tau T}}\log_{2}(1+\rho 2W\hat{p}_{\text{folded}}(f))df.}$ (13)
@@ V-A Signaling rate
One of the key benefits of FTN signaling is its ability to increase the number of parallel channels for a given TBP. To utilize the maximum number of signaling dimensions, the time-acceleration factor $\tau$ can be set to yield $N^{*}$ channels as follows:
@@
We see from (20) that $\tau^{*}$ approaches $\tau_{0}$ in the asymptotic TBP regime (i.e., as $\Omega\to\infty$). In the finite TBP regime, however, $\eta$ is typically smaller than $c$ for practical ranges of $\epsilon_{W}$, which implies $\tau^{*}<\tau_{0}$ is required to utilize all $N^{*}$ dimensions. The minimum $c$ needed to meet a given OOB power constraint can be inferred from Fig. 5, which plots the OOB energy versus $c$ for various pulses. For instance, $c\geq 10.52$ for the RRC pulse with $\beta=0.5$ when $\epsilon_{W}=10^{-4}$.
Fig. 6 plots the normalized SNRs of the parallel channels ($\sigma_{n}^{-2}/\rho$) for FTN systems with various $\tau$ values. As shown in solid lines, using $\tau_{0}$ activates only a partial number of channels, especially in low-TBP regimes, indicating that there is a benefit of selecting $\tau<\tau_{0}$ in the finite TBP regime. Setting $\tau=\tau^{*}$ yields $N^{*}$ channels as also indicated in the figure. The levels of these additional channels scale linearly with $\rho$ and thus the rate gains are higher in the high SNR. There is a diminishing return on rate gains for $\tau<\tau^{*}$, as the additional channels beyond $N^{*}$ possess diminishing SNRs and thus contribute little to the overall rate.
@@ V-B Time-limited pulse design
Various pulse designs have been proposed for FTN signaling, ranging from classical waveforms like the RRC, Gaussian, and PSWFs to pulses optimized for specific performance metrics. Liveris and Georghiades [liveris2003exploiting] established the existence of the Mazo limit for RRC pulses, demonstrating that they can be packed beyond the Nyquist rate without reducing the minimum Euclidean distance. Le et al. [le2014practical] evaluated several $T$-orthogonal pulses including RRC, Phydyas, and extended Gaussian functions regarding their bit error rate, peak-to-average power ratio, and spectral efficiency. Rusek et al. [rusek2008faster] explored short, finite pulses such as the half-cycle sinusoid and Halpern’s finite Nyquist pulse. However, the lack of a common OOB energy baseline makes direct comparisons difficult. The use of PSWFs for FTN appears in the early work of Foschini [foschini1984contrasting] and Anderson and Rusek [anderson2007optimal]. Notably, PSWFs and Gaussian pulses were identified to provide superior performance to RRC pulses when packet length is short [anderson2007optimal].
More recent literature focuses on optimized pulse design under some performance metrics. Anderson [anderson2017bandwidth, Ch. 7] explored several optimized pulses, including a modified principal PSWF, that simultaneously satisfy OOB and OOI requirements; however, these idealized pulses have infinite time and frequency support. Jaffal and Alvarado [jaffal2022pulses] proposed a weighted combination of normalized, time-truncated PSWFs, where weights are optimized to minimize residual ISI power (specifically power beyond the equalizer’s memory) under OOB constraints. Similarly, Jaffal and Abou-Faycal [jaffal2019achievable] utilized linear combination of PSWFs to maximize capacity in bandlimited channels. Diverging from orthogonal designs, Milojković et al. [milojkovic2022pulseshaping] introduced non-orthogonal pulses formed by linear combination of RRC pulses, optimized to maximize an information rate upper bound. Additionally, Makarov et al. [makarov2020optimizing] proposed time-limited pulses based on Fourier-series (FS) expansion, optimized for minimum OOB power subject to specific autocorrelation constraints. Despite these advancements, the time-truncated RRC pulse remains the common choice in the FTN literature due to its versatility and wide adoption in communication standards.
@@
We now compare the performance of the optimized pulse against other common pulses. Fig. 7 plots $R_{\text{NA}}$ for FTN systems using various pulses, alongside the PSWF benchmarks as detailed in Appendix E. The considered pulses include our optimized pulse from Table I (labeled ”modified Makarov”), time-truncated RRC pulses with various $\beta$, Anderson’s modified PSWF [anderson2017bandwidth], the principal PSWF pulse $\phi_{c,0}(t)$ (labeled ”PSWF0”), Gaussian pulse, and half-cycle sinusoid raised to $n$-th power (labeled $\sin^{n}$”). [^7] Anderson’s pulse is given by $p(t)=\alpha_{1}\psi_{c,0}(t)\text{ for }|t|\leq T/2$ and $p(t)=\alpha_{2}\psi_{c,0}(t)\text{ for }|t|>T/2$, where $\alpha_{1}$ and $\alpha_{2}$ are chosen so that OOB and OOI are both equal to $\epsilon$. See [anderson2017bandwidth] for details. Anderson’s pulse is further time-truncated to $|t|<T_{p}/2$ and energy normalized in our work. Gauss is a time-truncated and energy normalized Gaussian function with the parameter $\sigma=Q^{-1}\left(\epsilon_{W}/2\right)/(2\sqrt{2}\pi W)$, chosen to make OOB equal to $\epsilon_{W}$. Half-cycle sinusoid ($\sin^{n}$) is given by $\cos^{n}(\pi t/T_{p})$ for $|t|\leq T_{p}/2$ and $n\in\mathbb{Z^{+}}$, which is further normalized to have unit energy. For each pulse, the minimum pulse-width (or the minimum $c=2WT_{p}$) needed to meet the OOB constraint $\epsilon_{W}=10^{-4}$ is obtained from a plot of OOB energy versus $c$ in Fig. 5.888We use the conventional definition of $T=\tfrac{1+\beta}{2W}$ for RRC. The non-Nyquist pulses {Anderson, $\text{PSWF}_{0}$, Gauss, and $\text{sin}^{n}$}, however, do not have a well-defined Nyquist symbol time. For these pulses, we use the generalized definition of Nyquist symbol time proposed in [zhou2012generalized]. Specifically, the normalized bandwidth is first defined as $2B_{\text{norm}}=\tfrac{\int_{-W}^{W}|\hat{p}(f)|^{2}df}{\max_{f}|\hat{p}(f)|^{2}}$, which is the base-width of a rectangle whose height equals the peak of the pulse spectrum and whose area equals the in-band energy. The generalized Nyquist symbol time is then defined as $T=\tfrac{1}{2B_{\text{norm}}}$. The roll-off factor is also generalized as $\beta=\tfrac{W-B_{\text{norm}}}{B_{\text{norm}}}$. These definitions reduce to the conventional definitions when applied to RRC and sinc pulses.
@@ V-C Coding design
The FTN receiver implements three-stage turbo equalization between the FTN channel equalizer, URC decoder, and RSC decoder that iteratively exchange the extrinsic information of the coded bits. Instead of the low-complexity frequency-domain FTN equalizer considered in [sugiura2015frequency], we use the optimal maximum a posteriori (MAP) equalizer with $L$-memory that is operating on the trellis description of the FTN channel model (9). Leveraging on the fact that the FTN channel model is an instance of the Ungerboeck observation model, the MAP equalizer is implemented using the BCJR algorithm with modified branch metrics (see eq. (11) in [rusek2015ungerboeck]) [^9] The computational complexity of the optimal MAP equalizer is $\mathcal{O}(M^{L}N)$ for $L$-memory equalizer with $M$ modulation order and $N$ symbols. The exponential complexity can be replaced by polynomial (or even linear) complexity using reduced-complexity FTN equalizers at some performance loss [sugiura2015frequency], [bedeer2017very]. Fortunately, $L$ generally need not be large, since RRC pulses with $\beta$ close to 1 that is shown to be near-optimal in low TBP has fast temporal roll off and generates short ISI even when $\tau\ll 1$. As TBP increases, larger $\beta$ is needed but at the same time $\tau$ need not be small (both $\tau^{*}$ and $\tau_{0}$ get closer to 1) thus ISI remains short. Channel shortening technique [fan2018mlse, li2020code] can be used to further lower $L$ if needed. Numerical simulations confirmed that $L=3$ was sufficient to perform near MCCR in most cases of interest.. The MAP equalizer and the URC decoder form an inner turbo loop and exchange the extrinsic information about the URC encoded bits. The URC decoder generates the extrinsic information of both the URC encoded bits and the RSC encoded bits. After the inner loop iterates for $I_{\text{in}}$ times, the outer turbo loop between the URC decoder and the RSC decoder is run, in which the extrinsic information about the RSC encoded bits is exchanged. The inner loop then restarts with the updated extrinsic information from the RSC decoder becoming the a prior information about the URC encoded bits. The outer loop iterates for $I_{\text{out}}$, with which the total number of turbo iteration is $I_{\text{in}}I_{\text{out}}$. See [sugiura2015frequency] for more details.
@@ VI Conclusion
From a design perspective, we identified the optimal time-acceleration factor needed to maximize the available signaling dimensions and showed that with optimized pulse shaping, FTN’s performance can closely approach the theoretical benchmark set by PSWFs. Furthermore, we validated these theoretical gains with a practical three-stage turbo-equalization coding scheme that performs close to the MCCR bound, even for a TBP as low as $\Omega=132$. Further performance gain is expected by using more advanced and recent FTN receiver designs (e.g., [li2020code], [yang20256g]) by increasing the computational complexity.
Collectively, these results confirm that FTN effectively reduces the performance penalty from finite-length coding by creating more signaling dimensions within a constrained TBP. The fundamental benefits identified are expected to extend to more complex environments, such as fading, MIMO, and multi-user channels, positioning FTN as a valuable technique for future communication systems [li2026faster].
@@ Appendix A Proof of Theorem 2
$i_{\text{M}}(\tilde{x}_{n};\tilde{y}_{n})\triangleq\ln\left(\frac{p_{\tilde{y}|\tilde{x}}(\tilde{y}_{n}|\tilde{x}_{n})}{q_{\tilde{y}}(\tilde{y}_{n})}\right).$
@@
$i_{\text{M}}(\tilde{x}_{n};\tilde{y}_{n})=\ln\left(1+\frac{1}{\sigma_{n}^{2}}\right)-\frac{|\tilde{y}_{n}-(1+\sigma_{n}^{2})\tilde{x}_{n}|^{2}}{\sigma_{n}^{2}(1+\sigma_{n}^{2})}+|\tilde{x}_{n}|^{2}.$ (21)
@@ Appendix B Asymptotic expansion of the MC bound and proof of Proposition 1
$P_{\text{MD}}(\gamma)=P\left[\sum_{n=0}^{N-1}i_{\text{M}}(\tilde{x}_{n};\tilde{y}_{n})\leq N\gamma\middle|\text{H}_{1},\tilde{\mathbf{x}}\right$],
where $i_{\text{M}}(\tilde{x}_{n};\tilde{y}_{n})$ is the mismatched information density and given by (21) for the FTN channel. We will need its first three moments, which are given in the lemma below.
@@
(moments of mismatched information density): Let $|\tilde{x}_{n}|=1$. The mean and the variance of the mismatched information density $i_{\text{M}}(\tilde{x}_{n};\tilde{y}_{n})$ under $\tilde{y}_{n}\sim p_{\tilde{y}\mid\tilde{x}}=\mathcal{CN}(\tilde{x}_{n},\sigma_{n}^{2})$ and for a fixed $\tilde{x}_{n}$ are
$\begin{split}&c_{n}\triangleq\mathbb{E}_{\tilde{y}_{n}}[i_{\text{M}}(\tilde{x}_{n};\tilde{y}_{n})]=\ln\left(1+\tfrac{1}{\sigma_{n}^{2}}\right),\\ &v_{n}\triangleq\text{var}_{\tilde{y}_{n}}[i_{\text{M}}(\tilde{x}_{n};\tilde{y}_{n})]=1-\frac{1}{(1+\frac{1}{\sigma_{n}^{2}})^{2}}.\end{split}$
The third absolute central moment, $\theta_{n}\triangleq\mathbb{E}_{\tilde{y}_{n}}[|i_{\text{M}}(\tilde{x}_{n};\tilde{y}_{n})-c_{n}|^{3}]$, can be bounded as
@@
$(\mathbb{E}_{\tilde{y}_{n}}[|i_{\text{M}}(\tilde{x}_{n};\tilde{y}_{n})-c_{n}|^{2}])^{\frac{1}{2}}\leq\theta_{n}^{\frac{1}{3}}\leq(\mathbb{E}_{\tilde{y}_{n}}[|i_{\text{M}}(\tilde{x}_{n};\tilde{y}_{n})-c_{n}|^{4}])^{\frac{1}{4}}.$